# Germany, https://ais.uni-bonn.de

### v1

@@ 5 Conclusion
In this paper, we introduced LIAM, a multimodal model incorporating natural language input, ego-centric RGB image input, action history, and an accumulated semantic map. Because all the modality-specific encoders encode inputs in their own latent space, we designed two pre-training tasks to pre-align the embedding spaces. This pre-training objective aims to maximize the agreement of the action representation and the global representation of two consecutive images. However, the pre-training task does not consider the alignment of the text encoding backbone. We introduced a triple contrastive alignment to address this issue. The agreement between the image sequence representation and the text sequence representation, the two-frames representation and the action representation are expected to be maximized.
@@
[…]

### v2

@@ 5 Conclusion
In this paper, we introduced LIAM, a multimodal model incorporating natural language input, egocentric RGB image input, action history, and an accumulated semantic map. Because all the modality-specific encoders encode inputs in their own latent space, we designed two pre-training tasks to pre-align the embedding spaces. This pre-training objective aims to maximize the agreement of the action representation and the global representation of two consecutive images. However, the pre-training task does not consider the alignment of the text encoding backbone. We introduced a triple contrastive alignment to address this issue. The agreement between the image sequence representation and the text sequence representation, the two-frame representation and the action representation are expected to be maximized.
@@
Acknowledgements:
This work has been funded by the German Ministry of Education and Research under the grant no. 16SV8683, project: Transferzentrum Roboter im Alltag (RimA).